OpenStudy (ohmath123):

Assume f is the function defined by 2t+9 if t<0 3t-10 if t≥0 Evaluate f(1)

OpenStudy (trojanpoem):

1 > 0 3t- 10 t >= 0 ( plug here) f(1) = 3 * 1 - 10 = -7

OpenStudy (welshfella):

here t > 0 so use the second expression 3t - 10

OpenStudy (ohmath123):

im a little confused. how come you used the second expression? @TrojanPoem

OpenStudy (trojanpoem):

It asks to evaluate the expression when t = 1 f(1) the given intervals are 2t + 9 when t is smaller than zero 3t- 10 when t is greater than zero 1 is greater than zero so we pick the second as our f(t)

OpenStudy (ohmath123):

oh i see. Thank you!

OpenStudy (trojanpoem):

you're welcome.

OpenStudy (ohmath123):

so what if it now says evaluate f(|x|+1) @TrojanPoem ?

OpenStudy (trojanpoem):

f(x+1) x > 0 , f(-x+1) x < 0

OpenStudy (ohmath123):

i have to plug it in for both?

OpenStudy (trojanpoem):

yeah